The overall distance traveled by a golf ball is tested by hitting the ball with Iron Byron, a mechanical golfer with a swingPlease show all work and circle final answers. Thank you!
Answer #1

for brand I n-10 Ex = 2769 medio, sé, = 276.9 Standard deviation, si (7:23)2 W 8.7990 For brand 2 me =10 Ex: 2658 mean, ry= 2Test statistic t=ło-zła Sý limitlina) - 276.9. 265.8 92.5831 (0+110) - 2.7313 = 2.731 Degree of freedom dfan, t02-2 = 10 +10-b) Coritical value eft with significance level, 2=0.05 and df=18 is tdie,df=t0.0512, 18 = 2.1009 95% confidence interval CI =